OpenStudy (anonymous):

2+log2(2x−2)=2log2(x−1) the (log2),the 2 should be a bit down the log

jhonyy9 (jhonyy9):

2 is the base of log ?

OpenStudy (anonymous):

\[2+\log_{2} (2x-2)=2\log_{2}(x-1) \]

OpenStudy (anonymous):

right?

jhonyy9 (jhonyy9):

4(2x-2)=(x-1)(x-1)

OpenStudy (anonymous):

start with \[ 2=\log_2((x-1)^2)-\log_2(2(x-1))\]then \[2=\log_2(\frac{(x-1)^2}{2(x-1)})\] then \[2=\log_2(\frac{x-1}{2})\] and finally \[\frac{x-1}{2}=2^2=4\] and solve for x

OpenStudy (anonymous):

actually jhonny method is snappier!

OpenStudy (anonymous):

but still get \[x-1=8\] \[x=9\]

OpenStudy (anonymous):

thanks i got 9 as my answer too